OpenStudy (lukecrayonz):

4^(3x-2)=1/2^(2x) 4^(3x-2)=1/2^(2x) @Mathematics

OpenStudy (anonymous):

write both bases as powers of two \[4=2^2,\frac{1}{2}=2^{-1}\]

OpenStudy (lukecrayonz):

Ooh yesss :P Forgot about that. I got it from there

OpenStudy (anonymous):

so you get \[2^{2(3x-2)}=2^{-1(2x)}\] and then \[2(3x-2)=-2x\] etc

OpenStudy (anonymous):

got it

OpenStudy (lukecrayonz):

Thanks :D